Home Products common stock sells for $18.31 a share and has a market rate of return of 12.8 percent. The company just paid an annual dividend of $1.42 per share. What is the dividend growth rate?
Answer #1

Market rate of return=(Dividend payable next year/Stock price currently)+growth rate

Let the growth be g.

Hence dividend payable next year =(1.42*(1+g))

=$1.42+1.42g

Hence

0.128=[(1.42+1.42g)/18.31]+g

0.128=(1.42+1.42g+18.31g)/18.31

0.128*18.31=1.42+19.73g

Hence g=(2.34368-1.42)/19.73

=0.0468 (approx)

=4.68%(approx) is the dividend grwoth rate.
Answer #2

ANSWER :


Given :

r = 12.8% = 0.128 

P0 = 18.31 ($)

D0 = 1.42 ($)


Let dividend growth rate be g.. 

So, 

D1 = 1.42(1 + g) 

Now,

r = D/ P 0 + g

=> 0.128 = 1.42(1 + g) / 18.31 + g 

=> 0.128 = 1.42/18.31 + g (1.42/18.31 + 1) 

=> 0.128 = 0.07756 + 1.07756 g

=> g = (0.128 - 0.07756) / 1.07756

=> g = 0.04681

=> g = 4.681% = 4.68 % approx.


So,


4.68 % is the dividend growth rate. (ANSWER)


answered by: Tulsiram Garg